# CoCoGaussian: Leveraging Circle of Confusion for Gaussian Splatting from Defocused Images

### v1

@@ 1 Introduction
Recently, there has been increasing interest in reconstructing 3D scenes using only degraded images, with a focus on addressing challenges caused by camera motion blur and defocus blur. Deblur-NeRF [18] is the first study to take on this task, adopting a blind image deblurring approaches [40, 2, 31]. Follow-up studies [11, 25, 26, 13, 12, 37, 10] have expanded this approach beyond ray-tracing to rasterization-based methods. However, many of these methods depend heavily on learning-based strategies and fail to incorporate photographic principles that accurately capture how defocus occurs in real-world scenarios.
@@ 2.2 Novel View Synthesis from Blurry Images
Scene representation methods like NeRF [20] and 3DGS [9] require sharp images as input to render photo-realistic images. However, in real-world scenarios, capturing sharp images is challenging, often suffering from image degradation such as camera motion blur and defocus blur. To address this issue, Deblur-NeRF [18] introduced a 3D ray-based blurring kernel inspired by image blind deblurring [40, 2, 31], which intentionally generates blurry images during training. After training, it renders sharp novel-view images by excluding the trained kernel. Following Deblur-NeRF, various approaches have been proposed to tackle the degradation issue. DP-NeRF [11] introduces a kernel based on the rigid body transformation [17] with prior knowledge that object shape remains consistent in static scenes, and PDRF [25] proposes a blur estimation method with 2-stage efficient rendering. Recently, methods utilizing 3DGS have been proposed to enable faster rendering. For instance, Deblurring 3DGS [10] adjusts Gaussian parameters like rotation and scaling to generate blurry images during training, and BAGS [26] proposes a blur-agnostic kernel and a blur mask using convolutional neural networks (CNNs). In this paper, we propose a method that applies the photographic principles of defocus, utilizing only defocused images to physically model the circle of confusion (CoC).
@@ 3 Preliminary
Image blind deblurring is a technique for learning an image blurring kernel without the supervision of sharp images. Blurry images are obtained by applying convolution with a learned blurring kernel to sharp images, where the kernel size is fixed as a grid structure at each pixel position. Inspired by this approach, Deblur-NeRF [18] applies this technique to NeRF, presenting a method to represent a clean 3D scene using only blurry images. Specifically, Deblur-NeRF warps the input ray into multiple rays that constitute the blur, modeling a ray-based adaptive sparse kernel. A blurry pixel is obtained by combining the pixel colors of each ray:
@@ 5.1 Rendering Results
We quantitatively evaluate our CoCoGaussian using the following three metrics: peak signal-to-noise ratio (PSNR), structural similarity index measure (SSIM) [38], and learned perceptual image patch similarity (LPIPS) [45]. We compare our approach to several state-of-the-art methods, including both ray-tracing [18, 11, 25] and rasterization-based methods [26, 10]. Quantitative results are shown in Tab. 1 and Tab. 2, where our approach achieves the best performance across all metrics except for PSNR on the Deblur-NeRF real-world dataset. Note that we report the reproduced performance scores for Deblurring 3DGS [10] and BAGS [26] using their official codes. The Deblur-NeRF real-world dataset inherently presents challenges, as there are illumination differences between the sharp and defocused images, leading to relatively lower PSNR score. We discuss this issue in the appendix. For the DoF-NeRF dataset, CoCoGaussian achieves outstanding performance across all metrics, with an approximately 0.04 reduction in LPIPS compared to BAGS. Since the DoF-NeRF dataset has a higher resolution than the Deblur-NeRF dataset, this performance drop in BAGS is likely due to its requirement for manual adjustment of blurring kernel size, highlighting an intrinsic limitation. Conversely, Deblurring 3DGS shows minimal improvement over 3DGS, which can also be attributed to its limited kernel size relative to the higher resolution.
@@ 13 Per-Scene Quantitative Results
We present the performance for individual scenes across all datasets in Tabs. 6, 7 and 8. CoCoGaussian achieves the best LPIPS scores in all scenes except for the Coral scene in the Deblur-NeRF Real-World dataset. As discussed in Secs. 11 and 10, this indicates that CoCoGaussian exhibits the most superior high-level feature representation.

### v2

@@ Abstract
00footnotetext: This work was done during an internship at NAVER Cloud.
@@ 1 Introduction
Recently, there has been increasing interest in reconstructing 3D scenes using only degraded images, with a focus on addressing challenges caused by camera motion blur and defocus blur. Deblur-NeRF [18] is the first study to take on this task, adopting a blind image deblurring approaches [40, 2, 31]. Follow-up studies [11, 25, 26, 13, 12, 37, 10] have expanded this approach beyond ray-tracing to rasterization-based methods. However, many of these methods depend heavily on learning-based strategies and fail to incorporate photographic principles that accurately capture how defocus occurs in real-world scenarios. While DoF-NeRF [41] adopts the concept of CoC to model defocus blur in 3D scene representations, it relies on an implicit representation, which leads to slow rendering speeds, making it less suitable for real-time applications.
@@ 2.2 Novel View Synthesis from Blurry Images
Scene representation methods like NeRF [20] and 3DGS [9] require sharp images as input to render photo-realistic images. However, in real-world scenarios, capturing sharp images is challenging, often suffering from image degradation such as camera motion blur and defocus blur. To address this issue, Deblur-NeRF [18] introduced a 3D ray-based blurring kernel inspired by image blind deblurring [40, 2, 31], which intentionally generates blurry images during training. After training, it renders sharp novel-view images by excluding the trained kernel. Following Deblur-NeRF, various approaches have been proposed to tackle the degradation issue. DP-NeRF [11] introduces a kernel based on the rigid body transformation [17] with prior knowledge that object shape remains consistent in static scenes, and PDRF [25] proposes a blur estimation method with 2-stage efficient rendering. Recently, methods utilizing 3DGS have been proposed to enable faster rendering. For instance, Deblurring 3DGS [10] adjusts Gaussian parameters to generate blurry images during training, and BAGS [26] proposes a blur-agnostic kernel and a blur mask using convolutional neural networks (CNNs). DoF-NeRF [41] has explored modeling defocus blur by leveraging the circle of confusion (CoC) within an implicit neural representation. In this paper, we propose a method that also leverages CoC to model defocus, which enables real-time rendering, while still maintaining high-quality depth-of-field effects.
@@ 3 Preliminary
Image blind deblurring is a technique for learning an image blurring kernel without the supervision of sharp images. Blurry images are obtained by applying convolution with a learned blurring kernel to sharp images, where the kernel size is fixed as a grid structure at each pixel position. Inspired by this approach, Deblur-NeRF [18] applies this technique to NeRF, presenting a method to represent a clean 3D scene using only blurry images. Specifically, Deblur-NeRF warps the input ray into multiple rays that constitute the blur, modeling a ray-based sparse kernel. A blurry pixel is obtained by combining the pixel colors of each ray:
@@ 5.1 Rendering Results
We quantitatively evaluate our CoCoGaussian using the following three metrics: peak signal-to-noise ratio (PSNR), structural similarity index measure (SSIM) [38], and learned perceptual image patch similarity (LPIPS) [46]. We compare our approach to several state-of-the-art methods, including both ray-tracing [18, 11, 25] and rasterization-based methods [26, 10]. Quantitative results are shown in Tab. 1 and Tab. 2, where our approach achieves the best performance across all metrics except for PSNR on the Deblur-NeRF real-world dataset. Note that we report the reproduced performance scores for Deblurring 3DGS [10] and BAGS [26] using their official codes. The Deblur-NeRF real-world dataset inherently presents challenges, as there are illumination differences between the sharp and defocused images, leading to relatively lower PSNR score. We discuss this issue in the appendix. For the DoF-NeRF dataset, CoCoGaussian achieves outstanding performance across all metrics, with an approximately 0.04 reduction in LPIPS compared to BAGS. Since the DoF-NeRF dataset has a higher resolution than the Deblur-NeRF dataset, this performance drop in BAGS is likely due to its requirement for manual adjustment of blurring kernel size, highlighting an intrinsic limitation. Conversely, Deblurring 3DGS shows minimal improvement over 3DGS, which can also be attributed to its limited kernel size relative to the higher resolution. The details regarding the time complexity of our model are provided in the appendix.
@@ 7 Conclusion
Acknowledgements.
This project was supported by the NAVER Cloud Corporation and the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2024-00340745).
@@ 12 CoC Visualization
Table 6: Computational Efficiency and Speed. * indicates that the rendering speed is identical to that of the corresponding model.
Methods  GPU Memory  Training Time  Rendering Speed
(GB) (min)
BAGS [26] 4.7 47 *Mip-Splatting [45]
CoCoGaussian 5.3 45 *3DGS [9]
13 Computational Efficiency and Speed
We compare our GPU usage, training time, and rendering speed with BAGS [26], a state-of-the-art method, on the Deblur-NeRF real-world dataset using an NVIDIA RTX 3090. As shown in the Tab. 6, CoCoGaussian achieves comparable resource consumption and training time while delivering superior performance.
After the training phase, CoCoGaussian renders sharp images using only the $\mathbf{G}_{B}$ through a naive 3DGS. In the other words, the rendering speed and memory cost are the same as 3DGS alone. Therefore, our method is feasible for real-time applications that rely on 3DGS rendering speeds.
@@ 14 Per-Scene Quantitative Results
We present the performance for individual scenes across all datasets in Tabs. 7, 8 and 9. CoCoGaussian achieves the best LPIPS scores in all scenes except for the Coral scene in the Deblur-NeRF Real-World dataset. As discussed in Secs. 11 and 10, this indicates that CoCoGaussian exhibits the most superior high-level feature representation.